### v1

@@ A.5 Deriving the ATE Estimand
We leverage the causal graphical model constructed in the previous supplementary methods section to find the average treatment effect (ATE) of exposure $E$ on the vaccine uptake rate $\dot{V}$. The CGM of Supplementary Figure S4 shows a path $E_{t-1}\rightarrow\dot{\alpha}_{t-1}\rightarrow\alpha_{t}\rightarrow\dot{V}_{t}$, indicating that there is a causal chain of effects from $E$ to $\dot{V}$. In addition, there is a causal chain from $\alpha_{t-1}$ to both $\dot{V}$ and $E$, indicating that $\alpha_{t-1}$ is a confounding variable that must be considered. To account for these relationships, we use do-calculus pearlDoCalculusRevisited2012 , as implemented in the DoWhy python package, to find the generic form of the ATE:
$\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}[\dot{V_{t}}\mid\alpha_{t-1}]\Big{)}$
@@
To write this expression in terms of our model variables and parameters, we plug in the expression for $\dot{V_{t}}$ in terms of $E_{t-1}$ according to our model equations. We can look at the causal path through the CGM ($E_{t-1}\rightarrow\dot{\alpha}_{t-1}\rightarrow\alpha_{t-1}\rightarrow\dot{V}_{t}$) to find the relevant model equations:
@@
$\alpha_{t}\approx\dot{\alpha}_{t-1}\Delta t+\alpha_{t-1}$
$\dot{V}_{t}=\nu S_{t-1}(1-\alpha_{t-1}).$
Plugging in these expressions and $\Delta t=1$ for a one-day change, we get:
@@
Next, we need to take the expectation value over our data set and the joint posterior distribution of our parameters. Specifically, we compute an expectation value over our counties $i$, times $t$, and posterior samples $s$, and weight the expectation by county population, $N_{i}$. So the expression for $\mathbb{E}_{t,i,s}[\dot{V}_{t}]$ is rewritten as:
$\mathbb{E}_{t,i,s}[\dot{V}_{t}]\approx\frac{1}{(n_{t}-1)n_{s}(\sum_{i}N_{i})}\sum_{t=1}^{n_{t}}\sum_{i}\sum_{s}\\ N_{i}\nu_{s}S_{t-1,i}(1-((\gamma_{p,s}+\gamma_{e,s}E_{t-1,i})(1-\alpha_{t-1,i,s})+\alpha_{t-1,i,s}))$
@@
where $n_{t}$ and $n_{s}$ are the number of dates and posterior samples, respectively. Note that this quantity depends explicitly on $\alpha_{t-1}$, so, $\mathbb{E}[\dot{V}_{t}\mid\alpha_{t-1}]=\mathbb{E}_{t,i,s}[\dot{V}_{t}]$. Taking the derivative with respect to $E_{t-1}$ and collapsing the expectation value to a more concise notation, we get our estimand:
$\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}[\dot{V_{t}}\mid\alpha_{t-1}]\Big{)}$ $\approx-\frac{1}{(n_{t}-1)n_{s}(\sum_{i}N_{i})}\sum_{t=1}^{n_{t}}\sum_{i}\sum_{s}N_{i}\nu_{s}S_{t-1,i}\gamma_{e,s}(1-\alpha_{t-1,i,s})$
$=-\mathbb{E}_{t,i,s}\Big{[}\nu_{s}S_{t-1,i}\gamma_{e,s}(1-\alpha_{t-1,i,s})\Big{]}$
$=-\mathbb{E}\Big{[}\nu S_{t-1}\gamma_{e}(1-\alpha_{t-1})\Big{]}.$
@@
We can estimate the change in vaccine uptake in each county on each date using a linear approximation of the total effect. This approximation is valid under the assumption that the total change in vaccination is small. Below we justify this assumption by estimating the change with another, independent method. Using this approximation, we can write the total change in vaccination as
${\Delta V}_{t,i}$ $\approx E_{t-1,i}\Big{(}\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}[\dot{V_{t}}\mid\alpha_{t-1}]\Big{)}\Big{)}$
@@
and the total change in vaccine uptake over the whole time period and nation-wide population as
${\Delta V}=(N\Delta t)(\mathbb{E}_{t,i}[{\Delta V}_{t,i}])=N\Delta t\Big{(}\frac{1}{n_{t}(\sum_{i}N_{i})}\sum_{i,t}N_{i}{\Delta V}_{t,i}\Big{)}$
@@
where $N$ is the total population (about 337 million people in the US) and $\Delta t$ is the length of the observation period in days (192 days from February to August 2021). This results in a figure of about 750,000 vaccines prevented between February and August 2021 in the US.
@@
We also computed this quantity in an alternative way by simulating the SIRVA model with the inferred parameters in a counterfactual scenario where the exposure in all counties was set to zero. Comparing the real data with this scenario allowed us to infer that exposure accounted for 980,000 fewer vaccinations. This alternative estimate is comparable to that obtained with the ATE-based method and helps validate our assumption about the small effect size. However, the ATE-based method is better able to explicitly account for potential confounding factors.
@@ A.6 Estimating Case and Death Rates among the Unvaccinated Population
We can get a real value of this quantity by plugging in mean estimates of these probabilities for the cases that occurred over a particular time period (e.g. the previous day) and the number of vaccinations at that point in time:
@@
Computing these risks for an unvaccinated person in the United States over the period February-August 2021 using CDC data for cases, deaths, and vaccinations and values of $\lambda_{C}=0.93$ and $\lambda_{D}=0.94$ cdc_covid_vaccine_effectiveness ; Bernal2021_covid_efficacy , we find $P(C\mid\bar{V})=0.0387$ and $P(D\mid\bar{V})=0.00057$. This equates to 3,870 cases and 57 deaths per 100,000 unvaccinated people, which we use in the main text to estimate the numbers of cases and deaths resulting from exposure to antivaccine content on Twitter.

### v2

@@ A.5 Deriving the ATE Estimand
We leverage the causal graphical model constructed in the previous supplementary methods section to find the average treatment effect (ATE) of exposure per capita $E$ on the vaccine uptake rate per capita $\dot{V_{i}}/N_{i}$. The CGM of Supplementary Figure S4 shows a path $E_{t-1}\rightarrow\dot{\alpha}_{t-1}\rightarrow\alpha_{t}\rightarrow\dot{V}_{t}$, indicating that there is a causal chain of effects from $E$ to $\dot{V}$. In addition, there is a causal chain from $\alpha_{t-1}$ to both $\dot{V}$ and $E$, indicating that $\alpha_{t-1}$ is a confounding variable that must be considered. To account for these relationships, we use do-calculus [14], as implemented in the DoWhy python package, to find the generic form of the ATE:
$\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}\Big{[}\frac{\dot{V}_{t,i}}{N_{i}}\mid\alpha_{t-1}\Big{]}\Big{)}$
@@
To write this expression in terms of our model variables and parameters, we plug in the expression for $\dot{V_{t}}$ in terms of $E_{t-1}$ according to our model equations. We can look at the causal path through the CGM ($E_{t-1}\rightarrow\dot{\alpha}_{t-1}\rightarrow\alpha_{t-1}\rightarrow\dot{V}_{t}$) to find the relevant model equations (Eqs. 5, 6, 7):
$\dot{V}_{t}=\nu S_{t-1}(1-\alpha_{t-1})$
@@
$\alpha_{t}\approx\dot{\alpha}_{t-1}\Delta t+\alpha_{t-1}.$
Plugging in these expressions and $\Delta t=1$ for a one time step change, we get:
@@
Next, we need to take the expectation value over our data set and the joint posterior distribution of our parameters. Specifically, we compute an expectation value over our counties $i$, times $t$, and posterior samples $s$, and weight the expectation by county population, $N_{i}$. So the expression for $\mathbb{E}_{t,i,s}[\frac{\dot{V}_{t,i}}{N_{i}}]$ is rewritten as:
$\mathbb{E}_{t,i,s}\Big{[}\frac{\dot{V}_{t,i}}{N_{i}}\Big{]}\approx\frac{1}{(n_{t}-1)n_{s}(\sum_{i}N_{i})}\\ \sum_{t=1}^{n_{t}}\sum_{i}\sum_{s}N_{i}\frac{(\nu_{s}S_{t-1,i}(1-((\gamma_{p,s}+\gamma_{e,s}E_{t-1,i})(1-\alpha_{t-1,i,s})+\alpha_{t-1,i,s})))}{N_{i}}$
where $n_{t}$ and $n_{s}$ are the number of dates and posterior samples, respectively. Note that this quantity depends explicitly on $\alpha_{t-1}$, so, $\mathbb{E}[\frac{\dot{V}_{t,i}}{N_{i}}\mid\alpha_{t-1}]=\mathbb{E}_{t,i,s}[\frac{\dot{V}_{t,i}}{N_{i}}]$. Note also that the factor of $\frac{N_{i}}{N_{i}}=1$ results from the population-weighted averaging and our use of per-capita vaccination rates.
@@
Taking the derivative with respect to $E_{t-1}$ and collapsing the expectation value to a more concise notation, we get our estimand:
$\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}\Big{[}\frac{\dot{V}_{t,i}}{N_{i}}\mid\alpha_{t-1}\Big{]}\Big{)}$ $\approx-\frac{1}{(n_{t}-1)n_{s}(\sum_{i}N_{i})}\sum_{t=1}^{n_{t}}\sum_{i}\sum_{s}\nu_{s}S_{t-1,i}\gamma_{e,s}(1-\alpha_{t-1,i,s})$
$=-\mathbb{E}_{t,i,s}\Big{[}\frac{1}{N_{i}}\nu_{s}S_{t-1,i}\gamma_{e,s}(1-\alpha_{t-1,i,s})\Big{]}$
$=-\mathbb{E}\Big{[}\frac{1}{N_{i}}\nu S_{t-1}\gamma_{e}(1-\alpha_{t-1})\Big{]}.$ (11)
Note that in the second line of the above equation, we divide by $N_{i}$ because our expectation operation $\mathbb{E}$ is weighted by $N_{i}$.
@@
We can estimate the change in vaccine uptake in each county on each date using a linear approximation of the total effect. This approximation is valid under the assumption that the total change in vaccination is small. (Below we justify this assumption by estimating the change with another, independent method.) Using this approximation, we can write the total change in vaccination per day given exposure as
$\frac{\Delta\dot{V}_{t,i}}{N_{i}}$ $\approx E_{t-1,i}\Big{(}\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}_{t,i}\Big{[}\frac{\dot{V}_{t,i}}{N_{i}}\Big{]}\Big{)}\Big{)}$ (12)
@@
and the total change in vaccine uptake over the whole time period ($\Delta t$) and nation-wide population ($N$) as
${\Delta V}=N\Delta t\Big{(}\mathbb{E}_{t,i}\Big{[}\frac{{\Delta\dot{V}}_{t,i}}{N_{i}}\Big{]}\Big{)}.$ (13)
Now plugging in Eqs. 11 and 12 into Eq. 13, we get our final result:
$\Delta V$ $=N\Delta t~\mathbb{E}_{t,i}\Big{[}E_{t-1,i}\Big{(}\frac{d}{dE_{t-1}}\Big{(}\mathbb{E}_{t,i}\Big{[}\frac{\dot{V}_{t,i}}{N_{i}}\mid\alpha_{t-1}\Big{]}\Big{)}\Big{)}\Big{]}$
$=-N\Delta t~\mathbb{E}_{t,i}\Big{[}E_{t-1,i}\Big{(}\mathbb{E}_{t,i}\Big{[}\frac{1}{N_{i}}\nu S_{t-1}\gamma_{e}(1-\alpha_{t-1})\Big{]}\Big{)}\Big{]}$ (14)
@@
where $N$ is the total population (about 337 million people in the US), $\Delta t$ is the length of the observation period in days (192 days from February to August 2021), and all expectation operations $\mathbb{E}$ are weighted by populations $N_{i}$. This results in a figure of 14,086 vaccinations prevented between February and August 2021 in the US.
@@
We also computed this quantity in an alternative way by simulating the SIRVA model with the inferred parameters in a counterfactual scenario where the exposure in all counties was set to zero. Comparing the real data with this scenario allowed us to infer that exposure accounted for 980,000 fewer vaccinations. This alternative estimate is compatible with that obtained with the ATE-based method and helps validate our assumption about the small effect size. However, the ATE-based method is better able to explicitly account for potential confounding factors.
@@ A.6 Estimating Case and Death Rates among the Unvaccinated Population
We can get a real value of this quantity by plugging in mean estimates of these probabilities for the cases that occurred over a particular time period (e.g., the previous day) and the number of vaccinations at that point in time:
@@
Computing these risks for an unvaccinated person in the United States over the period February-August 2021 using CDC data for cases, deaths, and vaccinations and values of $\lambda_{C}=0.93$ and $\lambda_{D}=0.94$ [31, 32], we find $P(C\mid\bar{V})=0.0387$ and $P(D\mid\bar{V})=0.00057$. This equates to 511 cases and 8 deaths per 100,000 unvaccinated people, which we use in the main text to estimate the numbers of cases and deaths resulting from exposure to antivaccine content on Twitter.